# Ephemeral Rollups are All you Need

### v3

@@ A.1 EVM frameworks
One of the earliest on-chain gaming frameworks is MUD [mud]. Lattice, the company behind the MUD framework, was the first to introduce the ECS (Entity-Component-System) paradigm on-chain. While currently on Optimism, MUD aims to resolve the EVM limitation by introducing an opinionated database that indexes the game state from the chain and can be easily queried like any normal relational DB. This enables quicker access to the table of components that constitute an Autonomous World instance compared to retrieving the state directly from the chain. The focus of MUD is on modularity, featuring a system of libraries and plugin-ins. The biggest limitation is the ability to scale or add customization at the runtime level. Assuming the projected path to scale games built on MUD is with L3, this would fragment the state of the game deployed across new rollups. A similar OP-stack-derived engine is Keystone, developed by Curio, a rollup framework improving the performance of the EVM. Curio is focusing on the more responsive types of games, thanks to the introduction of ”ticks”, atomic units of time, that ultimately makes the game feel and look more responsive. Similarly to MUD, every game built on Keystone will have to deploy their own independent rollup, with the same fragmentation problem of an isolated state.
@@ A.4 Considerations
In their own unique ways, every framework is trying to increase the performance in terms of blocktime and throughput but introducing composability and network effects tradeoffs. Isolated app chains and horizontal scaling in its current form feel extremely short-sighted as it fragments the state and reduce the synergies with other protocols at the base layer. If it’s true that Autonomous Worlds introduce a new level of significance in games and enable virtual experiences that act as avenues for commerce, play and work, we believe blockchains settling large volume of real-world economic activity, RWO, consumer-facing applications, decentralized Physical infrastructure, and synchronizing state at low latency globally are a better fit for these kind of new experiences.
@@ Appendix B Scalability Approaches
A prevalent approach to address scalability challenges involves layers 2, protocols built on top of the base blockchain (Layer 1) to increase transaction speed and reduce costs, and partitioning traffic into isolated segments, known as sharding. Both methods share similarities and scales by segregating the state and computation into independent layers/shards. A similar technique is often employed in multiplayer games, where the multiple server instances manage traffic, e.g., sharding by geographic zones, whether virtual or real-world regions. Layer and Shards must be interoperable; for instance, a player might transition from one map area to another, effectively moving between shards. This method of scalability is incorporated in most game engine described in section LABEL:s:engines(directly or indirectly). This solution, taken to the extreme, can lead to app-specific L3s, where state and traffic are sharded for each app/game, an emerging scaling strategy for games built on the dojo and mud engines.

### v4

@@ A.1 EVM frameworks
One of the earliest on-chain gaming frameworks is MUD [mud]. Lattice, the company behind the MUD framework, was the first to introduce the ECS (Entity-Component-System) paradigm on-chain. While currently on Optimism, MUD aims to resolve the EVM limitation by introducing an opinionated database that indexes the game state from the chain and can be easily queried like any normal relational DB. This enables quicker access to the table of components that constitute an Autonomous World instance compared to retrieving the state directly from the chain. The focus of MUD is on modularity, featuring a system of libraries and plugin-ins. The biggest limitation is the ability to scale or add customization at the runtime level. If MUD scales games using L3, it would fragment the game state across new rollups. A similar OP-stack-derived engine is Keystone, developed by Curio, a rollup framework improving the performance of the EVM. Curio is focusing on the more responsive types of games, thanks to the introduction of ”ticks”, atomic units of time, that ultimately makes the game feel and look more responsive. Similarly to MUD, every game built on Keystone will have to deploy their own independent rollup, with the same fragmentation problem of an isolated state.
@@ A.4 Considerations
In their own unique ways, every framework is trying to increase the performance in terms of blocktime and throughput but introducing composability and network effects tradeoffs. Isolated app chains and current horizontal scaling approaches seem short-sighted, as they fragment the state and reduce synergies with other base layer protocols. If it’s true that Autonomous Worlds introduce a new level of significance in games and enable virtual experiences that act as avenues for commerce, play and work, we believe blockchains settling large volume of real-world economic activity, RWO, consumer-facing applications, decentralized Physical infrastructure, and synchronizing state at low latency globally are a better fit for these kind of new experiences.
@@ Appendix B Scalability Approaches
A prevalent approach to address scalability challenges involves layers 2, protocols built on top of the base blockchain (Layer 1) to increase transaction speed and reduce costs, and partitioning traffic into isolated segments, known as sharding. Both methods share similarities and scales by segregating the state and computation into independent layers/shards. A similar technique is often employed in multiplayer games, where the multiple server instances manage traffic, e.g., sharding by geographic zones, whether virtual or real-world regions. Layer and Shards must be interoperable; for instance, a player might transition from one map area to another, effectively moving between shards. This method of scalability is incorporated in most game engine described in section A(directly or indirectly). This solution, taken to the extreme, can lead to app-specific L3s, where state and traffic are sharded for each app/game, an emerging scaling strategy for games built on the dojo and mud engines.